OpenStudy (anonymous):

I have the auxiliary equation \[r^2-4r+13=0\] The roots are imaginary, but I forgot how to determine them.

OpenStudy (anonymous):

complete the square is easiest here \[x^2-4x=-13\] \[(x-2)^2=-13+4=-9\] \[x-2=\pm\sqrt{-9}=\pm3i\] \[x=2\pm3i\]

OpenStudy (anonymous):

makes sense. Thank you!

OpenStudy (anonymous):

yw